# ARIM-mdx Data System: Towards a Nationwide Data Platform for Materials Science

### v2

@@ I Introduction
II Background: ARIM Japan Project
Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM Japan 2021–2031) [29] is a continuation of a series of long-term initiatives (Nanotechnology Network Project in 2007–2011, Low Carbon Research Network Project in 2011–2016, Nanotech Technology Platform Project in 2012–2021), that aim to effectively manage shared-use experimental facilities located broadly in 13 universities and research institutes. Over 1,000 units of experimental equipment are registered for shared use, and more than 2,000 research projects utilize them annually.
@@ Acknowledgment
The authors would like to thank all the contributors [^2] Contributors https://arim.mdx.jp/en/#devteam. We used mdx [47] and the HPCI/JHPCN supercomputers [^3] Supercomputers https://arim.mdx.jp/en/#supercomputer. This work was supported by MEXT ARIM Japan, “Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN)” (jh241010), MEXT “Developing a Research Data Ecosystem for the Promotion of Data-Driven Science” (2024-10) and JSPS KAKENHI (22K17899).

### v3

@@ I Introduction
II Background: MEXT ARIM Japan Project
Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM Japan 2021–2031) [29] by the Ministry of Education, Culture, Sports, Science, and Technology (MEXT) is a continuation of a series of long-term initiatives (Nanotechnology Network Project in 2007–2011, Low Carbon Research Network Project in 2011–2016, Nanotech Technology Platform Project in 2012–2021), that aim to effectively manage shared-use experimental facilities located broadly in 13 universities and research institutes. Over 1,000 units of experimental equipment are registered for shared use, and more than 2,000 research projects utilize them annually.
@@ Acknowledgment
The authors would like to thank all the contributors [^2] Contributors https://arim.mdx.jp/en/#devteam. We used mdx [47] and the HPCI/JHPCN supercomputers [^3] Supercomputers https://arim.mdx.jp/en/#supercomputer. This work was supported by MEXT ARIM Japan, MEXT “Data Creation and Utilisation Type Materials Research and Development Project (DxMT)” (JPMXP1121467561), “Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN)” (jh241010), MEXT “Developing a Research Data Ecosystem for the Promotion of Data-Driven Science” (2024-10) and JSPS KAKENHI (22K17899). We used ChatGPT 4o and Claude 3.5 Sonnet for grammar checks and writing revisions.